OpenStudy (anonymous):

Can someone please help me... Simplify completely: 10x^6y^3+20x^3y^2 ______________________ 5x^3y

OpenStudy (anonymous):

Can you use the equation button so I can see it clearer?

OpenStudy (hoblos):

10x^3y^2( x^3y + 2) = ---------------------- 5x^3y = 2y(x^3y +2) = 2x^3y^2 + 4y

OpenStudy (anonymous):